$A$ woman has $10$ keys,out of which only one opens a lock. She tries the keys one after another (keeping aside the failed ones) until she succeeds in opening the lock. What is the probability that the seventh key works?

  • A
    $\frac{7}{10}$
  • B
    $\frac{1}{2}$
  • C
    $\frac{3}{10}$
  • D
    $\frac{1}{10}$

Explore More

Similar Questions

When a missile is fired from a ship,the probability that it is intercepted is $\frac{1}{3}$ and the probability that the missile hits the target,given that it is not intercepted,is $\frac{3}{4}$. If three missiles are fired independently from the ship,then the probability that all three hit the target is:

Three cards are drawn successively,without replacement from a pack of $52$ well-shuffled cards. What is the probability that the first two cards are kings and the third card drawn is an ace?

There are $3$ bags $A, B$ and $C$. Bag $A$ contains $2$ white and $3$ black balls,bag $B$ contains $4$ white and $2$ black balls and Bag $C$ contains $3$ white and $2$ black balls. If a ball is drawn at random from a randomly chosen bag,then the probability that the ball drawn is black,is

Two balls are drawn at random with replacement from a box containing $10$ black and $8$ red balls. Find the probability that the first ball is black and the second is red.

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo